OpenStudy (khantahmina):

determine all point on the circle x^2+y^2=25 where the slope is 3/4

OpenStudy (steve816):

First step: Implicitely differentiate the equation!

OpenStudy (steve816):

I know you can do this. I trust you.

OpenStudy (khantahmina):

hehe thanks!!!! 2x+2y=0

OpenStudy (khantahmina):

ooh nvm

OpenStudy (khantahmina):

\[2x+2y(dy/dx)=0\]

OpenStudy (khantahmina):

dy/dx=-x/y

OpenStudy (khantahmina):

i went on and found dy/dx

OpenStudy (khantahmina):

is that ok

OpenStudy (steve816):

Yes, that is fine.

OpenStudy (khantahmina):

so then do i do \[3/4=-x/y\]

OpenStudy (khantahmina):

then idk what to dooooo

OpenStudy (steve816):

You then need to solve for y in the original equation.

OpenStudy (steve816):

And plug that in for y for the derivative.

OpenStudy (khantahmina):

i got \[y=\sqrt(25-x^2)\]

OpenStudy (steve816):

It would be + or -

OpenStudy (steve816):

Now, plug that in to your derivative equation. And finally, solve!

OpenStudy (khantahmina):

\[dy/dx= -x/(\sqrt(25-x^2))\]

OpenStudy (khantahmina):

right

OpenStudy (khantahmina):

?

OpenStudy (steve816):

Yeah, and remember that dy/dx is 3/4

OpenStudy (khantahmina):

i got x=3 x=-3

OpenStudy (steve816):

That looks about right! Congratulations!!!

OpenStudy (khantahmina):

thats it?

OpenStudy (sshayer):

\[x=-\frac{ 3 }{ 4 }y\] \[x^2+y^2=25\] \[\frac{ 9 }{ 16 }y^2+y^2=25\] \[9y^2+16y^2=25\] \[25 y^2=25\] \[y^2=1,y=\pm 1\] find x for each.

OpenStudy (sshayer):

correction

OpenStudy (steve816):

@sshayer Are you sure you are trying to find the point when the tangent line is -3/4? I don't see where you are getting at.

OpenStudy (sshayer):

\[9 y^2+16y^2=25*16=400,25 y^2=400,y^2=\frac{ 400 }{ 25 }=16,y=\pm 4\] find x for each.

OpenStudy (sshayer):

remember x and y are of different sign